the junior class sold 120 turkey dinner plates and 200 chicken dinner plates for a total of $2,150. the senior class sold 100 turkey plates and 300 chicken dinner plates for a total of $2,625. i need to know how much the turkey plates costs. that's all they gave us was the equation
Answer by mananth(16949)

Junior Class
120 turkey dinner plates
200 chicken dinner plates
total of $2,150.
....
the senior class
100 turkey plates
300 chicken dinner plates
total of $2,625.
..
Let turkey dinner plate cost $x
Chicken dinner plates cost $y
...
Junior ---120x+200y = 2150..................1
Senior ----100x+300y = 2625.................2
multiply (1) by -3 & (2) by 2
-360x-600y=-6450
200x+600y = 5250
..
Add both equations
-160x= -1200
/-160
x= -1200/-160
x=7.50
$ 7.50 for Turkey dinner plate
..
m.ananth@hotmail.ca